# Sample Motion for Structured Illumination Fluorescence Microscopy

### v1

@@ 5 Discussion
Sample motion between captured images has traditionally been viewed as a problem for super-resolution microscopy, causing motion blurring artifacts. Here, we have utilized it as a novel contrast mechanism for SIM. Rather than relying on time-varying excitation patterns, we rely on sample motion to encode diverse information into the measurements even with a fixed excitation pattern. This greatly simplifies the implementation of SIM systems. With a fixed pattern, optomechanical components or spatial light modulators are no longer needed to update the excitation patterns, consequently reducing the equipment cost as well as eliminating the time delay for the pattern update. Furthermore, a fixed speckle pattern is also easier to pre-calibrate compared to previous speckle SIM methods [2], which enhances the data efficiency of the reconstruction.
A limitation of our work is that the motion contrast mechanism may not encode sufficient information when some parts of a sample stay static. Then, the raw data will encode identical information over multiple exposures, resulting in an ill-posed inverse problem. This problem can be resolved by introducing external global motion to the sample, like shifting the sample laterally as in Fig. 5.
@@
In addition, fast sample motion may also blur individual raw images. Since NSTM assumes the sample remains motionless during each exposure, any movement within a single exposure cannot be accounted for. This can lead to information loss during encoding and cause model mismatch in motion estimation. Therefore, it is essential to carefully select the exposure settings to minimize blur in the raw images, which is a common requirement in other microscopy methods.
In summary, we introduced Speckle Flow SIM, which employs a static speckle illumination pattern to encode super-resolution information in fluorescence imaging. We demonstrated its feasibility through simulation analysis and proof-of-concept experiments. In our experiments, Speckle Flow SIM achieved super-resolution reconstruction on a continuously moving fluorescent-labeled sample. This motion contrast mechanism creates new opportunities for innovative optical designs for SIM systems.

### v2

@@ 5 Discussion
Sample motion between captured images has traditionally been viewed as a problem for super-resolution microscopy, causing motion blurring artifacts. Here, we have utilized it as a novel contrast mechanism for SIM. Rather than relying on time-varying excitation patterns, we rely on sample motion to encode diverse information into the measurements with a fixed excitation pattern. This greatly simplifies the implementation of SIM systems and removes motion blurring artifacts. With a fixed pattern, optomechanical components or spatial light modulators are no longer needed to update the excitation patterns, consequently reducing the equipment cost as well as eliminating the time delay for the pattern update. Furthermore, a single fixed speckle pattern is easy to pre-calibrate, unlike traditional speckle SIM, which enhances the data efficiency of the reconstruction.
A limitation of our work is that it requires motion to achieve super-resolution. For static parts of the sample, the raw data will encode identical information over all exposures, resulting in an ill-posed inverse problem. This can be mitigated by introducing external global motion, like shifting the sample laterally.
@@
If motion is very fast, then individual raw images may become blurred. Such motion blurring is not accounted for by our current algorithm; it could lead to information loss during encoding and cause model mismatch in motion estimation. Therefore, it is essential to carefully select the exposure settings to minimize blur in the raw images, which is a common requirement in other microscopy methods. Since the NSTM can predict motion during each raw image’s exposure time, future work could attempt to deblur the raw images computationally.
In summary, Speckle Flow SIM employs a static speckle illumination pattern to encode super-resolution information in fluorescence imaging. We demonstrated its feasibility through simulation analysis and proof-of-concept experiments. In our experiments, Speckle Flow SIM achieved super-resolution reconstruction on a continuously moving fluorescent-labeled sample. This motion contrast mechanism creates new opportunities for innovative optical designs for SIM systems.
[…]
Funding Chan Zuckerberg Initiative (DAF2021-225666, DAF2021-225666 (DOI: 10.37921/192752jrgbnh)); Chan Zuckerberg Initiative (Biohub SF investigator (L.W.)); National Science Foundation (DMR 1548924); Siebel Scholars Foundation (Siebel Scholar (R.C.)).
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability Statement Data and code in this paper are available in https://github.com/rmcao/SpeckleFlowSIM-fluo .